Given the area in the first quadrant bounded by x² = 8y, the line



x=4 and the x-axis. What is the



the volume generated by revolving this



area about the y-axis?

1
Expert's answer
2022-06-07T11:56:35-0400




"4^2=8y=>y=2"

"x^2=8y, x\ge0=>x=\sqrt{8y}"

"V=\pi \displaystyle\int_{0}^{2}(4^2-(\sqrt{8y})^2)dy"

"=\pi[16y-4y^2]\begin{matrix}
 2\\
 0
\end{matrix}=\pi(32-16)"

"=16\pi ({units}^3)"